A particle moves in space so that at time t its position is stated as x=2t+3, y=t2+3t, z=t3+2t2 Find the components of its velocity and acceleration when t=1


1
Expert's answer
2021-08-17T10:39:28-0400

Let a particle moves in space so that at time "t" its position is stated as "x=2t+3, y=t^2+3t, z=t^3+2t^2."

Let us find the components of its velocity:

"v_x=x'=2, v_y=y'=2t+3, v_z=z'=3t^2+4t."

Let us find the components of its acceleration:

"a_x=x''=0, a_y=y''=2, a_z=z''=6t+4."

When "t=1" the components of its velocity are "v_x=2, v_y=5, v_z=7" and the components of its acceleration are "a_x=0, a_y=2, a_z=10."